7<bo v 6 st J4 lum 06 6 Co use nodal Analysis to find vo and un
Answer #1

ti 7 268 Vn VO {gor --08 GLO because Comes under Super Node Technique * This question two soa Nodes & Un. Voltage source blsubstitute un value in equation @ Yo= 7 (68 + Vm = 7L60 + 40.49 (44.66° y = 47.27 %46.91° volls Answers - Vo = 47.27 (46.91